# Chiral excitation flows of multinode network based on synthetic gauge fields

### v1

@@ IV Extension of four-node networks
where the relevant parameters are based on the four-node case. All of the hopping phases between nearest-neighboring nodes are set to be $\theta_{0}=\pi/2$ to ensure the chiral symmetry of the system, and the coupling strengths between auxiliary nodes and network nodes are the same as the coupling in the four-node case, i.e., $\beta_{c}=2$. However, this configuration cannot be interpreted as a synthetic uniform magnetic field model because the hopping phases in this structure cannot be the same within a uniform magnetic field. In practice, it can still be synthesized in superconducting circuits by accurately setting the Floquet modulation parameters [63], see Sec. VI.1 for details.
@@
Figure 6: (a) Illustration of the ladder model with auxiliary nodes, where the corner nodes are labeled as $c_{j},~{}j=1,2,3,4$. The solid and dashed arrows are hoppings partitioned into $H_{0}$ and $V_{I}$, respectively. (b) and (c) correspond to the chiral flows with different copies of the four-node network($N=2$ and $N=6$ respectively). (e) The effective configuration of the ladder network. In this pattern, the corner nodes are equivalent. (f) The fidelity of chiral flow in the ladder network as a function of $N$. Here, the blue solid line represents the fidelity from identical coupling $\beta_{c}=2$, while the red dashed line denotes the results from optimal parameters with $\beta_{j}>2$, $j=1,\cdots,m$.
@@
We study the chiral dynamics of this ladder network by initiating an excitation at the 1-st node. Figs. 6(b) and 6(c) display the evolution of excited-state populations of all nodes for $N=2$ and $N=6$ cases respectively, which demonstrate that the primary characteristic of the chiral flow is retained. The majority of the excitation is chirally transported between the corner nodes $c_{j}$ [as labeled in Fig. 6(a)], namely, $21\rightarrow N+1\rightarrow N+2\rightarrow 2N+$ 2. We emphasize that the imperfection of this chiral flow originates from the unevenly spaced oscillation frequencies, as detailed in Appendix D, where we also provide the analytic expression of $P_{e}^{1}(t)$ by partitioning the Hamiltonian as $H_{\mathrm{ex}}^{N}=H_{0}+V_{I}$ and implementing the nonperturbative calculation of system’s propagator. As the number of copies $N$ increases, the timescale for transferring an excitation from node $1$ to node $N+1$ also extends (in a sublinear form). To describe the transmission rate of this chiral flow, we define the fidelity as $\mathcal{F}=|\langle\psi(t=0)|\psi_{f}\rangle|^{2}$, where $\psi_{f}$ is the final state after a cycle period. As shown in Fig. 6(f), the fidelity of chiral flow $\mathcal{F}$ decays linearly when the network becomes larger, with less remainder excitation at the 1-st node.
@@
Nevertheless, this extended model can also be viewed as a polygonal network with auxiliary nodes, as shown in Fig. 6(d). Within this pattern, we can still maintain equal treatment of the corner nodes by replacing $\beta_{c}$ with a sequence of coupling parameters $\beta_{j},j=0,1,\cdots m$, where $m=\lceil N/2\rceil-1$. Thus, we can then optimize the network by modulating the parameters. Intuitively, the coupling $\beta_{j}J_{0}$ should be enhanced to connect further nodes when it becomes far away from the corner atoms, i.e., $2=\beta_{0}<\beta_{1}<\cdots<\beta_{m}$. As illustrated in Fig. 6(f), we plot the optimized fidelity as a function of $N$ with a set of optimal parameters, which shows a linear decline but is improved measurably compared to the identical coupling case. Even if the network expands to ten times the four-node network’s scale, the fidelity remains at $\mathcal{F}>0.8$. Therefore, we are convinced that the chirality persists and dominates within a larger network, rendering this ladder network highly practical for implementing chiral transmission between remote nodes.
@@ V Imperfections in networks
In practical systems, the inevitable imperfections from various aspects of quantum networks can affect the quality of chiral flows in the $n$-node aSGF network and ladder networks. To estimate such influence, in this section, we investigate the sensitivity of chiral flow to the disorder on the node (bosonic mode or emitter) frequency $\omega_{0}$, coupling strength $J_{0}$, and hopping phase $\theta_{ij}$, respectively. Here, we take the four-node aSGF model as an example for illustration.
As shown in Fig. 7(a), these three types of disorder are denoted as $\delta_{\square}$, $\square=\omega,J$ and $\theta$, where the first acts on the node frequency while the latter two affect the hoppings between nodes. In the single-excitation subspace, the disorder on frequency and hopping is represented as the diagonal and off-diagonal disorder of the single-particle Hamiltonian $\mathcal{H}_{\mathrm{aSGF}}^{4}$, respectively. Here, we first consider the disorder on the frequency by replacing $\omega_{0}$ with $\omega_{0}+\delta_{\omega}$ for each node, where $\delta_{\omega}$ is sampled from a uniform distribution within the range [$-\delta_{\omega}^{m},\delta_{\omega}^{m}$]. Since the maximum excitation population of $j$-th node is no longer 1, i.e., $\max\left[P_{e}^{j}(t)\right]<1$ within a typical cycle period, we use the average fidelity of the corner excitation to characterize the quality of chiral flow:
@@
As illustrated in Fig. 7(a), the frequency disorder significantly influences the chiral flow. Here, the parameters in this boson-SGF network are taken from the experimental data [23], with $\omega_{0}=5.6~{}\mathrm{GHz}$ and $J_{0}=2\pi\times 4.29~{}\mathrm{MHz}$. Compared with the hopping, even a slight deviation of the node frequency can greatly change the energy spectrum, leading to a considerable decrease of the fidelity $\mathcal{F}_{\mathrm{avg}}$. Moreover, the chiral symmetry is broken when such diagonal (i.e., frequency) disorder is introduced, which results in the asymmetry of the energy spectrum and thus violates the criteria for conducting a perfect chiral flow. In contrast, the off-diagonal disorder preserves the chiral symmetry of the system and causes minor impacts on the energy spectrum. As plotted in Figs. 7(c) and 7(d), the chiral flow shows robustness ($\mathcal{F}_{\mathrm{avg}}>0.9$ with about $30\%$ deviation of the parameters) when the disorder stems from the hopping strength and phase.
Figure 7: (a) Illustration of the four-node aSGF model with inevitable imperfections. The imperfections of this network are described by the disorder of the node frequencies and hoppings. (b)-(d) Effects of different types of disorder on the average fidelity.
@@ VI Feasibility in experiments
In this section, we elucidate the practical implementation of our proposal, where we take the four-node aSGF model as an example in the following discussion. Here, we briefly introduce the schemes for superconducting circuits [22, 23] and magnetic photonic lattice [11, 18, 12, 52] platforms to achieve controllable hoppings with phases, thus promising the realization of multinode chiral flows in bosonic and atomic systems, respectively.
@@ VI.1 Superconducting circuits
Following the first way, we can design a superconducting circuit consisting of five transmon qubits (or resonators) $Q_{1}$, $Q_{2}$, $Q_{3}$, $Q_{4}$, $Q_{a}$, coupled with each other via tunable couplers, where the couplings can be dynamically modulated as $g_{jk}(t)=2g_{jk}\cos\left(\nu_{jk}t+\phi_{jk}\right)$ [66]. Here, the modulation frequency $\nu_{jk}$ is chosen to be the frequency detuning between elements $j$ and $k$, i.e., $\nu_{jk}=\omega_{j}-\omega_{k}$, and we assume that $\nu_{jk}\gg g$. After transforming to the interaction picture under the rotating-wave approximation (RWA), we obtain the effective Hamiltonian
@@
where $H_{\mathrm{nl}}\approx-U/2\left(\sum_{j}a_{j}^{\dagger}a_{j}^{\dagger}a_{j}a_{j}+c^{\dagger}c^{\dagger}cc\right)$ is the Kerr nonlinearity and arises from the cosine potential of transmon qubit, with the on-site potential $U\gg g$. The nonlinearity of superconducting qubits naturally promotes the effective Hamiltonian as a spin-aSGF model since it prohibits the double-photon occupancy if the transmons are initially singly occupied with double excitation like $a_{j}^{\dagger}a_{k}^{\dagger}|G\rangle$. For harmonic resonators with $U=0$, $H_{\mathrm{eff}}$ mimics a boson-SGF Hamiltonian. To satisfy the requirements of conducting a four-node chiral flow, we set the parameters as follows: $g_{j}^{c}/2=g_{j,j+1\mathrm{mod}4}=g$ for $j=1$, 2, 3, 4, $g_{13}=g_{24}=0$, $\phi_{j}^{c}=0$ and $\phi_{j,j+1\mathrm{mod}4}=-\pi/2$. Notably, the frequency detunings of the elements have to satisfy the RWA conditions: $|\omega_{j}-\omega_{k}|>g$ for $j\neq k$. Then, the Hamiltonian reduces to the form in Eq. (3) with $N=4$, $J_{0}=g$, and $\beta_{c}=2$, allowing for conducting the perfect chiral flow.
In addition, the first line of Hamiltonian (20) can also be synthesized via a bus resonator [23]. In this circuit quantum electrodynamics (QED) model, four elements are interconnected by a bus resonator with couplings $g_{j}$. To construct synthetic gauge fields, we periodically adjust the frequencies of qubits around the frequency of the bus resonator $\omega_{r}$ as
@@
where $\Delta$ and $\nu$ represent the modulation amplitude and frequency, with $\nu\approx\Delta/2.40$, and the phases are denoted as $\phi_{j}=j\pi/2$. Under the condition $\nu\gg g$, we obtain the effective first-order Hamiltonian with complex hopping amplitudes $g_{jk}={g_{j}g_{k}}\beta_{jk}e^{i\pi/2}/{\nu}$, by neglecting higher-order terms. Here, $\beta_{jk}$ is given by the expression $\beta_{jk}=\sum_{n=1}^{+\infty}2J_{n}^{2}\left[f\right]\sin\left[n\left(\phi_{k}-\phi_{j}\right)\right]/n$, where $J_{n}[x]$ represents the $n$th-order Bessel function of the first kind and the dimensionless variable $f=\Delta/\nu\approx 2.40$ [68, 69] such that the zeroth-order term $J_{0}[f]=0$. The $\pi/2$ hopping phase of $g_{jk}$ thus can accurately replicate the four-node chiral flow Hamiltonian. By setting $\phi_{j}=j\pi/2$, we can establish NN hoppings with the same strength while preventing the occurrence of NNN hoppings between nodes $j$ and $j+2$.
VI.2 Emitters in magnetic photonic lattice
@@
The circuit QED provides a convenient way for synthesizing hopping phases between artificial atoms or resonators. Here, we show that a hybrid system, i.e., a magnetic photonic lattice (MPL) coupled to atomic emitters [52], can also induce hopping phases and be utilized to construct our model. Although the chiral flow between emitters is imperfect, we still see this dominant feature in such a hybrid system.
@@
As shown in Fig. 8(a), the MPL is a square array composed of identical photonic resonators with equal spacing $l_{0}$ and frequency $\omega_{r}$, where the neighboring lattices are coupled via a synthetic magnetic field $\mathbf{B}$ with complex tunneling amplitudes. We place five atomic emitters to the resonators on the center and four corners of a square with side length $d$. These emitters have the same resonant frequency $\omega_{e}$ and are labeled as $1,2,3,4$, and $a$. The coupling strengths between the five emitters and resonators are identical, denoted as $g$.
@@
Figure 8: (a) Schematic of an aSGF model in a magnetic photonic lattice. The blue hollow circles are resonators that form a square lattice, while the filled circles with five different colors are emitters coupled to the resonators locally. The dashed lines between emitters indicate the effective long-range dipole-dipole interactions induced by the magnetic photonic lattice. (b) Evolution of the excited-state populations of emitters in the MPL, where $\ell=0$, $g/(\omega_{e}-\omega_{\ell=0})=0.1$ and $d/l_{0}=4$. And the unit of time is $\Omega^{-1}=(d/l_{0})g^{-1}$.
In the continuum limit, the Hamiltonian of the MPL can be approximately diagonalized to a series of Landau levels with energy $\omega_{\ell}$, each of which possesses a large number of degenerate states. Here, we assume that the nearest Landau level to the frequencies of emitters is $\ell=0$. Under conditions $\omega_{c}\gg g$ and $|\omega_{e}-\omega_{\ell=0}|\gg g$, a Schrieffer-Wolff transformation can be applied to obtain the effective atomic interaction Hamiltonian

### v2

@@ IV Extension of four-node networks
where the relevant parameters are based on the four-node case. The index $j$ in the second term runs over the nodes belonging to copy $i$ of the four-node network. Each node, except the corners, belongs to two copies. This ladder network bears some resemblance to quasi-1D topological insulators such as the rhombus chain [73] and the Creutz ladder [74]. All of the hopping phases between nearest-neighboring nodes are set to be $\theta_{0}=\pi/2$ to ensure the chiral symmetry of the system, and the coupling strengths between auxiliary nodes and network nodes are the same as the coupling in the four-node case, i.e., $\beta_{c}=2$; however, this configuration cannot be interpreted as a synthetic uniform magnetic field model because the hopping phases in this structure cannot be the same within a uniform magnetic field. In practice, it can still be synthesized in superconducting circuits by accurately setting the Floquet modulation parameters [75] (see Sec. VI.1 for details).
@@
Figure 6: (a) Illustration of the ladder model with auxiliary nodes, where the corner nodes are labeled as $\mathscr{C}_{j}(j=1,2,3,4)$. The solid and dashed arrows are hoppings partitioned into $H_{0}$ and $V_{I}$, respectively. (b),(c) Chiral flows with different copies of the four-node network($N=2$ and $N=6$ respectively). (d) Effective configuration of the ladder network. In this pattern, the corner nodes are equivalent. (e) Fidelity of chiral flow in the ladder network as a function of $N$. Here, the blue solid line represents the fidelity from identical coupling $\beta_{c}=2$, while the red dashed line denotes the results from optimal parameters with $\beta_{j}>2$ ($j=1,\cdots,m$).
@@
We study the chiral dynamics of this ladder network by initiating an excitation at the first node. Figures 6(b) and 6(c) display the evolution of the excited-state populations of all nodes for the $N=2$ and $N=6$ cases respectively, demonstrating that the primary characteristic of the chiral flow is retained. The majority of the excitation is chirally transported between the corner nodes $\mathscr{C}_{j}$ [as labeled in Fig. 6(a)], namely, $21\rightarrow N+1\rightarrow N+2\rightarrow 2N+$ 2. We emphasize that the imperfection of this chiral flow originates from the unevenly spaced oscillation frequencies, as detailed in Appendix D, where we also provide the analytic expression for $P_{e}^{1}(t)$ by partitioning the Hamiltonian as $H_{\mathrm{ex}}^{N}=H_{0}+V_{I}$ and implementing the nonperturbative calculation of system’s propagator. As the number of copies $N$ increases, the timescale for transferring an excitation from node $1$ to node $N+1$ also increases in a linear form (see Appendix E). To describe the transmission rate of this chiral flow, we define the fidelity as $\mathcal{F}=|\langle\psi(t=0)|\psi_{f}\rangle|^{2}$, where $\psi_{f}$ is the final state after a cycle period. As shown in Fig. 6(e), the fidelity of chiral flow $\mathcal{F}$ decays linearly when the network becomes larger, with less remaining excitation at the first node.
@@
The extended four-node model can also be viewed as a polygonal network with auxiliary nodes, as shown in Fig. 6(d). Within this pattern, we can still maintain equal treatment of the corner nodes by replacing $\beta_{c}$ with a sequence of coupling parameters $\beta_{j}(j=0,1,\cdots m$, where $m=\lceil N/2\rceil-1$). Thus, we can then optimize the network by modulating the parameters. Intuitively, the coupling $\beta_{j}J_{0}$ should be enhanced to connect further nodes when it becomes distant from the corner nodes, i.e., $2=\beta_{0}<\beta_{1}<\cdots<\beta_{m}$. This has been verified by numerical calculation, where we systematically iterate through these parameters to identify the optimal values (see Appendix E). In Fig. 6(e), we plot the optimized fidelity as a function of $N$ with optimal parameters, which shows a linear decline but is improved measurably compared to the identical coupling case. Even if the network expands to ten times the four-node network’s scale, the fidelity remains at $\mathcal{F}>0.8$. Therefore, we are convinced that the chirality persists and dominates within a larger network, rendering this ladder network highly practical for implementing chiral transmission between remote nodes.
@@ V Imperfections in networks
In practical systems, the inevitable imperfections from various aspects of quantum networks can affect the quality of chiral flows in the $n$-node aSGF network and ladder networks. To estimate such influence, in this section, we investigate the sensitivity of chiral flow to the disorder on the node (bosonic mode or atomic transition) frequency $\omega_{0}$, coupling strength $J_{0}$, and hopping phase $\theta_{ij}$. Here, we take the four-node aSGF model as an example for illustration.
As shown in Fig. 7(a), these three types of disorder are denoted as $\delta_{\square}$, $\square=\omega,J$, and $\theta$, where the first acts on the node frequency while the latter two affect the hoppings between nodes. In the single-excitation subspace, the disorder on frequency and hopping are represented as the diagonal and off-diagonal disorder of the single-particle Hamiltonian $\mathcal{H}_{\mathrm{aSGF}}^{4}$, respectively. We first consider the disorder on the frequency by replacing $\omega_{0}$ with $\omega_{0}+\delta_{\omega}$ for each node, where $\delta_{\omega}$ is sampled from a uniform distribution within the range [$-\delta_{\omega}^{m},\delta_{\omega}^{m}$]. Similarly, the coupling strengths and hopping phases are sampled among $[J_{0}-\delta^{m}_{J},J_{0}+\delta^{m}_{J}]$ and $[\pi/2-\delta^{m}_{\theta},\pi/2+\delta^{m}_{\theta}]$, respectively. In our simulations, the parameters are taken from experimental data based on a superconducting circuit platform [30], where transmon qubits serve as artificial atoms and complex hopping terms are implemented via a superconducting bus resonator (see Sec. VI.1 for details). In such a platform, imperfections primarily arise from fabrication-induced disorder, unintended crosstalk between closely spaced circuit elements, and material defects, all of which collectively lead to deviations from the designed parameters. Specifically, we use $\omega_{0}=5.6~{}\mathrm{GHz}$, which represents the frequency of each artificial atom, and $J_{0}=2\pi\times 4.29~{}\mathrm{MHz}$, which denotes the hopping strength. Since the maximum excitation population of the $j$th node is no longer 1, i.e., $\max\left[P_{e}^{j}(t)\right]<1$ within a typical cycle period, we use the average fidelity of the corner excitation to characterize the quality of chiral flow:
@@
As illustrated in Fig. 7(b), the frequency disorder significantly influences the chiral flow. Compared with the hopping, even a slight deviation of the node frequency can greatly change the energy spectrum, leading to a considerable decrease in the fidelity $\mathcal{F}_{\mathrm{avg}}$. Moreover, the chiral symmetry is broken when such diagonal (i.e., frequency) disorder is introduced, which results in asymmetry of the energy spectrum and thus violates the criteria for conducting a perfect chiral flow. In contrast, off-diagonal disorder preserves the chiral symmetry of the system and has only minor impacts on the energy spectrum. As plotted in Figs. 7(c) and 7(d), the chiral flow shows robustness ($\mathcal{F}_{\mathrm{avg}}>0.9$ with about $30\%$ deviation of the parameters) when the disorder stems from the hopping strength and phase.
Figure 7: (a) Illustration of the four-node aSGF model with inevitable imperfections. The imperfections of this network are described by the disorder of the node frequencies and hoppings. (b)-(d) Effects of different types of disorder on the average fidelity. The numerical data represent the average values derived from 2000 instances of disorder. The error bars in the graph signify one standard deviation of the data points.
@@ VI Feasibility in experiments
In this section, we elucidate the practical implementation of our proposal, where we take the four-node aSGF model as an example in the discussion that follows. Here, we first briefly introduce the schemes for superconducting circuits [29, 30] and magnetic photonic lattice platforms [11, 18, 12, 59] to achieve controllable hoppings with phases, thus promising the realization of multinode chiral flows in bosonic and atomic systems, respectively. We then elaborate on the synthesis of gauge fields in ultracold atoms, which offers another promising platform for multinode chiral flows in both spin and bosonic systems.
@@ VI.1 Superconducting circuits
Following the first approach, we can design a superconducting circuit consisting of five transmon qubits (or resonators) $Q_{1}$, $Q_{2}$, $Q_{3}$, $Q_{4}$ and $Q_{a}$, coupled with each other via tunable couplers, where the couplings can be dynamically modulated as $g_{jk}(t)=2g_{jk}\cos\left(\nu_{jk}t+\phi_{jk}\right)$ [78]. Here, the modulation frequency $\nu_{jk}$ is chosen to be the frequency detuning between elements $j$ and $k$, i.e., $\nu_{jk}=\omega_{j}-\omega_{k}$, and we assume that $\nu_{jk}\gg g$. After transforming to the interaction picture under the rotating-wave approximation (RWA), we obtain the effective Hamiltonian
@@
where $H_{\mathrm{nl}}\approx-U/2\left(\sum_{j}a_{j}^{\dagger}a_{j}^{\dagger}a_{j}a_{j}+c^{\dagger}c^{\dagger}cc\right)$ is the Kerr nonlinearity and arises from the cosine potential of a transmon qubit, with the on-site potential $U\gg g$. The nonlinearity of superconducting qubits naturally promotes the effective Hamiltonian as a spin-aSGF model since it prohibits the double-photon occupancy if the transmons are initially singly occupied with double excitation denoted by $a_{j}^{\dagger}a_{k}^{\dagger}|G\rangle$. For harmonic resonators with $U=0$, $H_{\mathrm{eff}}$ mimics a boson-SGF Hamiltonian. To satisfy the requirements of conducting a four-node chiral flow, we set the parameters as follows: $g_{j}^{c}/2=g_{j,j+1\mathrm{mod}4}=g$ for $j=1$, 2, 3, and 4; $g_{13}=g_{24}=0$; $\phi_{j}^{c}=0$; and $\phi_{j,j+1\mathrm{mod}4}=-\pi/2$. Notably, the frequency detunings of the elements have to satisfy the RWA conditions: $|\omega_{j}-\omega_{k}|>g$ for $j\neq k$. The Hamiltonian then reduces to the form in Eq. (3) with $N=4$, $J_{0}=g$, and $\beta_{c}=2$, allowing for the conduction of the perfect chiral flow.
In addition, the first sum of Hamiltonian (20) can be synthesized via a bus resonator [30]. In this circuit quantum electrodynamics (QED) model, four elements are interconnected by a bus resonator with couplings $g_{j}$. To construct synthetic gauge fields, we periodically adjust the frequencies of qubits around the frequency of the bus resonator $\omega_{r}$ as
@@
where $\Delta$ and $\nu$ represent the modulation amplitude and frequency, with $\nu\approx\Delta/2.40$, and the phases are denoted as $\phi_{j}=j\pi/2$. Under the condition $\nu\gg g$, we obtain the effective first-order Hamiltonian with complex hopping amplitudes $g_{jk}={g_{j}g_{k}}\beta_{jk}e^{i\pi/2}/{\nu}$, by neglecting higher-order terms. Here, $\beta_{jk}$ is given by the expression $\beta_{jk}=\sum_{n=1}^{+\infty}2J_{n}^{2}\left[f\right]\sin\left[n\left(\phi_{k}-\phi_{j}\right)\right]/n$, where $J_{n}[x]$ represents the $n$th-order Bessel function of the first kind and the dimensionless variable $f=\Delta/\nu\approx 2.40$ [80, 81] such that the zeroth-order term $J_{0}[f]=0$. The $\pi/2$ hopping phase of $g_{jk}$ can thus accurately replicate the four-node chiral-flow Hamiltonian. By setting $\phi_{j}=j\pi/2$, we can establish NN hoppings with the same strength while preventing the occurrence of NNN hoppings between nodes $j$ and $j+2$.
VI.2 Emitters in magnetic photonic lattices
@@
The circuit QED provides a convenient way to synthesize hopping phases between artificial atoms or resonators. Here, we show that a hybrid system, i.e., a magnetic photonic lattice (MPL) coupled to atomic emitters [59], can also induce hopping phases and be used to construct our model. Although the chiral flow between emitters is imperfect, we still see this dominant feature in such a hybrid system.
@@
As shown in Fig. 8(a), the MPL is a square array composed of identical photonic resonators with equal spacing $l_{0}$ and frequency $\omega_{r}$, where the neighboring lattices are coupled via a synthetic magnetic field $\bm{B}$ perpendicular to the array. The Hamiltonian of the MPL is a tight-binding lattice model with complex tunneling amplitudes. It is given by
$H_{\mathrm{MPL}}=\omega_{r}\sum_{i}\Psi^{\dagger}_{i}\Psi_{i}-J\sum_{\langle i,j\rangle}(e^{i\theta_{ij}}\Psi^{\dagger}_{i}\Psi_{j}+\mathrm{H.c.}),$ (22)
where $\Psi^{(\dagger)}_{i}$ annihilates (creates) a photon of frequency $\omega_{r}$ inside the resonator at $\bm{r}_{i}$. In the continuum limit, this Hamiltonian can be approximately diagonalized into a series of Landau levels with energies $\omega_{\ell}\sim BJ(\ell+1/2)$ ($\ell=0,1,2,\dots$), each of which possesses a large number of degenerate states [59].
We place five atomic emitters in resonators located at the four corners and the center of a square with side length $d$. These two-level emitters are described by the Pauli matrices $\sigma^{j}$ and are labeled as 1, 2, 3, 4, and $a$. They have the same emitter frequency $\omega_{e}$. Each emitter is locally coupled to its corresponding on-site resonator through a standard Jaynes-Cummings term, $g(\Psi_{j}\sigma_{+}^{j}+\Psi^{\dagger}_{j}\sigma_{-}^{j})$, where $g$ represents the uniform coupling strength for all five emitter-resonator pairs. The MPL acts as a structured photonic bath, mediating photon-exchange interactions among the emitters through collective photonic modes. If the coupling strength is much smaller than the system’s Landau energy gap, i.e., $g\ll\omega_{\ell+1}-\omega_{\ell}$, the emitters are effectively coupled to the Landau level closest to $\omega_{e}$, which is assumed to be $\ell=0$. Under the condition of large detuning $|\omega_{e}-\omega_{\ell=0}|\gg g$, we can apply a Schrieffer-Wolff transformation to obtain the effective Hamiltonian
@@
Figure 8: (a) Schematic of an aSGF model in a magnetic photonic lattice. The blue hollow circles are resonators that form a square lattice, while the filled circles with five different colors are emitters coupled to the resonators locally. The dashed lines between emitters indicate the effective long-range dipole-dipole interactions induced by the magnetic photonic lattice. (b) Evolution of the excited-state populations of emitters in the MPL, where $\ell=0$, $g/(\omega_{e}-\omega_{\ell=0})=0.1$, the distance between nearest emitter is $|\bm{r}_{j}-\bm{r}_{j+1}|/l_{0}=4$ and the unit of time is $\Omega^{-1}=(|\bm{r}_{j}-\bm{r}_{j+1}|/l_{0})g^{-1}$.